40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Apr 8 08:05:41 2013, Steve, Update, PSL, PMC locked 
    Reply  Mon Apr 8 10:19:46 2013, Jenne, Update, PSL, LSC left enabled 
    Reply  Tue Apr 9 07:48:10 2013, Steve, Update, PSL, PMC locked pmc10d.png
       Reply  Wed Apr 10 07:37:33 2013, Steve, Update, PSL, PMC locked pmc4d.png
Message ID: 8422     Entry time: Mon Apr 8 10:19:46 2013     In reply to: 8421
Author: Jenne 
Type: Update 
Category: PSL 
Subject: LSC left enabled 

 

 Note: The TRY PD isn't installed and normalized properly yet, so the IFO OVERVIEW screen indicates lock for the Yarm constantly, which is not true.  Hopefully in the next day or so the screen will be back to telling the truth.

Also, the LSC Locking was left ENABLED (presumably over the weekend).   This is not so good.  It can kick optics around, so we should all take a look when we walk through the control room, and if no one is locking, please disable the LSC master switch. 

ELOG V3.1.3-